OpenStudy (anonymous):

A man has two types of coffee beans. The first type sells for $5.20/lb, the second for $5.80/lb. How many pounds of the first type must be mixed with 9lb of the second type to produce a blend selling for $5.35/lb?

OpenStudy (anonymous):

Hi there!. This is a system of linear equations.

OpenStudy (anonymous):

So what you want to do is set up an equation that computes the weighted-average cost of the two types of beans. Do you know what a weighted average is?

OpenStudy (anonymous):

No I am not so sure

OpenStudy (anonymous):

Thank you for helping me! :)

OpenStudy (anonymous):

OK. Well, you probably know what a regular average is, right?

OpenStudy (anonymous):

A regular average of two numbers is, for example, (x + y)/2, or (x + y + z)/3. The number in the denominator is the number of items we're averaging over.

OpenStudy (anonymous):

Imagine however, that we want an average over three things, but two are the same: x+x+y. The average is (x+x+y)/3, since we still have 3 items, even though two are the same. We can rewrite this as (2x + y) / 3 -- and now we have a weighted average. It's the sum of each value, times its weight, divided by the total of the weights. Does that make sense? It means x has twice as much effect on the result as y.

OpenStudy (anonymous):

Does that make sense so far?

OpenStudy (anonymous):

Yes that makes sense!

OpenStudy (anonymous):

Great! So, we have two kinds of beans, and they have two different costs. We know the weight -- in this case, literally a weight in pounds -- of one of the kinds, but not the other. We'll call the unknown amount "x". So the weighted average of the two is: \[{ { (x * 5.20) * (9 * 5.80) } \over {x+9} }= 5.35 \]

OpenStudy (anonymous):

x*5.20 is the value of the first bean, 9*5.80 is the value of the second bean, x+9 is the total weight of the beans. Divide them, and you get dollars per pound. We set that equal to what we *want* the dollars per pound to be, or 5.35. Does that make sense?

OpenStudy (anonymous):

Oops, that was supposed to be a plus -- sorry, let me write it out again. \[{ {(x*5.20) + (9*5.80) } \over x+9} = 5.35 \]

OpenStudy (anonymous):

Oh that's okay!

OpenStudy (anonymous):

In most of math, finding the right equation that represents your problem is the hard part. Solving the equation usually is the easy part. (Computers can usually do that for you, in fact)

OpenStudy (anonymous):

Okay so would I simply solve for the equation you typed?? I really appreciate your help!

OpenStudy (anonymous):

Yep, that's it! That will give you a value for x, which is the number of pounds of the other kind of bean. To solve the equation, a good first step is to multiply both sides by (x+9).

OpenStudy (anonymous):

that looks like a good method. we can also reason as follows: put x = pounds of first beans used, they cost \[5.20x\]and 9 pounds of the other beans will cost \[9\times 5.80=52.50\]for a total of \[5.20x+52.50\] if you add x pounds of beans to the 9 pounds of beans you will have \[x+9\] pounds of beans, and you want this to be worth \[5.35(x+52.50)\] so set them equal and solve that is write \[5.20x+52.50=5.35(x+9)\] and solve for x

OpenStudy (anonymous):

Thank you both so very much for your help!! Did you get an end result? On the first method I got around 3.65lbs but the second I got almost 30 lbs

OpenStudy (anonymous):

i get x = 29

OpenStudy (anonymous):

I get 27 pounds

OpenStudy (anonymous):

i could have made a mistake

OpenStudy (anonymous):

mathematica solved it as 27

OpenStudy (anonymous):

oh well let me try again

OpenStudy (anonymous):

I might have made a typo in the input though; here's what it says: In[819]:= Solve[(5.20 x + (9*5.8))/(x + 9) == 5.35, {x}] Out[819]= {{x -> 27.}}

OpenStudy (anonymous):

Oh that makes sense! I just retried it and got 27 as well.! :)

OpenStudy (anonymous):

if you start with \[520x+5250=535(x+9)\] i think we can do it easily

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

Thank you SO much for all of your time I really appreciate it!!!!

OpenStudy (anonymous):

I hope it was helpful :-)

OpenStudy (anonymous):

i still get 29 did i make a mistake in the formula?

OpenStudy (anonymous):

It definitely was! :)

OpenStudy (anonymous):

yeah, 9*5.8 is 5220, not 5250

OpenStudy (anonymous):

ahhhhhhhh so it is!